OpenStudy (anonymous):

write in standard form the equation of each line. 1) slope= 3; (-1,4) I know standard form is ax+by=c but I have no idea how to even begin solving this problem let alone understand the entire subject please help

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) ... start with point slope form y - y1 = 3(x - x1) ... plug in the given slope y - 4 = 3(x - (-1)) ... plug in the given point y - 4 = 3(x + 1) y - 4 = 3x + 3 y = 3x + 3 + 4 y = 3x + 7 y - 3x = 7 -3x + y = 7 3x - y = -7

jimthompson5910 (jim_thompson5910):

Let me know if that helps or not

OpenStudy (anonymous):

wait, so y=3x+7 how did you get y-3x=7? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I subtracted 3x from both sides to get y - 3x = 7 then rearranged the equation to get -3x + y = 7

jimthompson5910 (jim_thompson5910):

Most books will want a > 0, so I then multiplied every term in -3x + y = 7 by -1 to get 3x - y = -7

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

ok, so for 2) slope= -1; (0,7) y-y1=m(x-x1) y-y1=-1(x-0) y-7=-1x-0 y=7-1x 1x+y=7 Is that right?

jimthompson5910 (jim_thompson5910):

1x+y=7 is the same as x+y=7

jimthompson5910 (jim_thompson5910):

Let's check that answer x + y = 7 0 + 7 = 7 ... plug in x = 0, y = 7. This is from (0,7) 7 = 7 ... this is true, so (0,7) definitely lies on this line

jimthompson5910 (jim_thompson5910):

And it turns out that x+y = 7 has a slope of -1 because solving for y gives you y = -x + 7

jimthompson5910 (jim_thompson5910):

So the answer checks out

OpenStudy (anonymous):

I did it! oh my gosh thank you SO much, I was struggling with this subject

jimthompson5910 (jim_thompson5910):

Yes you did. Just keep practicing with this and it'll click much smoother. However, it looks that you are already a pro at this. So congratulations with that. Keep up the good work.

OpenStudy (anonymous):

thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome